# MAS-Bench: A Unified Benchmark for Shortcut-Augmented Hybrid Mobile GUI Agents

### v1

@@ 5.3 Result of Shortcut Generation
In contrast, Replay-Task shortcuts exhibit poor robustness with only a 10% success rate, indicating high susceptibility to environmental variations. This fragility negatively impacts both task completion rates and execution efficiency, underscoring that while effective shortcuts boost GUI performance, poorly designed ones are counterproductive.
Our findings reveal substantial room for improvement in model-generated shortcuts. Among all variants, only Predefined and MobileAgent-E shortcuts improve model accuracy, with only Predefined shortcuts simultaneously enhancing both execution speed and success rate. These results indicate current limitations in shortcut generation capabilities within our framework. Future research should focus on developing more efficient, robust shortcuts with higher utilization rates that effectively reduce execution steps.
Shortcut
[…]
@@ 6 Conclusion
In this paper, we introduce MAS-Bench, the first benchmark designed to evaluate the effectiveness of GUI-shortcut hybrid mobile agents. Furthermore, MAS-Bench provides a framework to access the quality and validity of shortcuts that are autonomously generated by the agent. Our experiments demonstrate that the GUI-shortcut hybrid operational approach enhances both the success rate and efficiency of task execution. Moreover, we also validate that our benchmark effectively measures the quality and robustness of the agent-generated shortcuts.
Limitations and Future Work. Despite our efforts to provide a comprehensive evaluation, the range of comparative baselines remains limited due to the currently underexplored research in Mobile GUI-shortcut hybrid Agents. Therefore, we hope to pay more attention to the efficiency of mobile GUI agents in the future, promoting research into GUI-shortcut Hybrid Mobile Agents.
7 Acknowledgments
This work was supported by Huzhou Natural Science Foundation under Grant 2024YZ01 and in part by the Postdoctoral Fellowship Program of CPSF under Grant GZC20241491.
@@ A.1 Observation Space
MAS-Bench evaluates mobile agents within a highly standardized and controllable environment using an Android Virtual Device (AVD). Specifically, the environment is based on a Pixel 7 emulator running Android 15 (API Level 35). Its screen resolution is set to 1080 x 2400 pixels, and the screen density is 420 dpi. The emulator has 24GB of RAM and 24GB of storage. The environment can support parallel execution of multi-threaded evaluations.
A.2 Details of Action Space
@@
A.3 Details of Predefined Shortcuts Knowledge Base
@@
Our shortcut knowledge base was created to provide agents with an authentic yet challenging environment. We collect and design shortcuts for 11 selected applications, ensuring their relevance to real-world usage patterns.
  • APIs Collection: The API shortcuts in our knowledge base are application-specific internal APIs. We identify these APIs through official Android documentation. We employ static analysis of the application packages for internal APIs that lack public documentation to identify key functional endpoints.
@@
  • RPAs Script Design: RPA scripts are designed for common, highly repetitive sub-tasks (e.g., adding an item with specific options to a shopping cart). These scripts can be executed as a single action, serving as a high-level shortcut. This design enables us to test an agent’s ability to utilize complex, pre-built automation routines.
@@
MobileAgentV2 Gemini-2.0-Flash 0.250 1.280 472.426 40.406 0 0
MAS-MobileAgent Gemini-2.0-Flash 0.402 0.719 335.181 39.597 3.565 0.505
MobileAgentV2 Gemini-2.5-Pro 0.446 1.058 1013.386 120.212 0 0
MAS-MobileAgent Gemini-2.5-Pro 0.641 0.613 682.547 99.780 1.348 0.345
@@
MobileAgentV2 Gemini-2.0-Flash 0 - 952.660 106.501 0 0
MAS-MobileAgent Gemini-2.0-Flash 0.234 0.738 403.798 71.350 5.362 0.709
MobileAgentV2 Gemini-2.5-Pro 0.170 1.247 2053.133 227.128 0 0
MAS-MobileAgent Gemini-2.5-Pro 0.617 0.829 1441.586 189.836 3.128 0.320
Table 5: Evaluation results of different base models. MSRS is the Mean Step Ratio on Successful tasks, MET is the Mean Execution Time in seconds, and MToC is the Mean Token Cost in thousands (kTokens).
A.4 Details of Tasks in MAS-Bench
@@
This suite comprises 139 complex tasks strategically designed across 11 popular real-world applications. The tasks are categorized into 92 single-app tasks and 47 cross-app workflows, with their difficulty distribution illustrated in Table 6. We have also curated a corresponding knowledge base of predefined shortcuts to support these tasks. Table 9 presents a breakdown of the selected applications and the number of single-app tasks and shortcuts associated with each.
@@ Appendix B Details of Predefined Shortcut Evaluation
B.1 Details of MAS-MobileAgent and MAS-T3A

### v2

@@ 5.4 Results of Shortcut Generation
In contrast, Replay-Task shortcuts exhibit poor robustness with only a 10% success rate, indicating high susceptibility to environmental variations. This fragility negatively impacts task completion rates and execution efficiency, underscoring that while effective shortcuts boost GUI performance, poorly designed ones are counterproductive.
Results reveal substantial room for improvement in model-generated shortcuts. Across variants, only Predefined and MobileAgent-E shortcuts improve model accuracy, with only Predefined shortcuts simultaneously enhancing execution speed and success rate. These results indicate current limitations in shortcut generation capabilities within our framework. Future work should focus on developing more efficient, robust shortcuts with higher utilization rates that effectively reduce execution steps.
@@ 6 Conclusion
In this paper, we introduce MAS-Bench, the first unified and comprehensive benchmark designed to evaluate the effectiveness of GUI-shortcut hybrid mobile agents. Furthermore, MAS-Bench provides a framework to assess the quality and validity of shortcuts that are autonomously generated by the agent. Our experiments demonstrate that the GUI-shortcut hybrid operation significantly enhances both the success rate and efficiency of task execution. Moreover, we also validate that our benchmark effectively measures the quality and robustness of the agent-generated shortcuts. We hope MAS-Bench will support future work on efficient GUI-shortcut hybrid mobile agents, including agent self-improvement through iterative shortcut accumulation, validation, and reuse.
7 Limitations
While our predefined shortcuts demonstrate robustness across application versions (Appendix A.4), agents should maintain the capability to fall back to GUI operations when shortcuts occasionally fail due to environmental variations. Additionally, our evaluation on standardized emulators ensures reproducibility and fair comparison; future work will validate these findings on real-world devices with diverse configurations. Another direction is to integrate dynamic shortcut discovery into MAS-Bench, e.g., by mining AndroidManifest files, static analysis results, official documentation, or web resources to construct structural shortcuts automatically. Finally, although MAS-Bench-Eval shows strong agreement with human judgments, future versions can combine rule-based state checks with LLM-as-a-Judge to further improve evaluation stability. This work highlights the importance of efficiency in mobile GUI agents and calls for further research into GUI-shortcut hybrid approaches.
Appendix
@@ A.1 Observation Space
MAS-Bench evaluates mobile agents in a highly standardized and controllable environment using an Android Virtual Device (AVD). Specifically, the environment is based on a Pixel 7 emulator running Android 15 (API Level 35). Its screen resolution is set to 1080 x 2400 pixels, and the screen density is 420 dpi. The emulator has 24GB of RAM and 24GB of storage. The environment support parallel execution of multi-threaded evaluations. All model deployments and evaluations are conducted on 4 NVIDIA L40S GPUs 48GB.
A.2 Environment Control and Reproducibility
MAS-Bench’s dynamic evaluation approach moves beyond simply verifying task outcomes, focusing on the agent’s decision-making process in live scenarios. To ensure real-world relevance, the benchmark incorporates a suite of widely used Android applications covering diverse daily-life scenarios. Ensuring consistent and reproducible evaluation environments is critical for fair agent comparison, particularly when working with dynamic online applications. MAS-Bench addresses this challenge through a snapshot-based rapid recovery mechanism combined with carefully controlled test accounts.
Snapshot-Based Rapid Recovery.
We leverage Android emulator snapshot functionality to guarantee absolute environment consistency across evaluations. Prior to experimentation, we construct a pre-configured Android Virtual Device image, termed MAS-Bench-AVD, which includes all 11 benchmark applications with necessary permissions granted and dedicated test accounts logged in. This clean baseline state is saved as a snapshot. We maintain a single system-level snapshot rather than task- or app-specific snapshots, so this design introduces no additional per-task storage overhead beyond the base AVD image.
Online Content Control.
For applications involving dynamic online content (e.g., Amazon, YouTube, Gmail), we employ pre-registered anonymous dedicated test accounts to maintain a relatively stable recommendation baseline. These accounts are isolated from real user data. For tasks involving email communication, we utilize anonymous test accounts configured with automatic reply functionality, ensuring that all email-based tasks interact only within our controlled test environment. This design prevents interference with real-world users while maintaining realistic task scenarios.
Parallel Execution and Scalability.
To support large-scale evaluation, MAS-Bench implements ADB port-based isolation for managing multiple emulator instances concurrently. Each emulator instance operates independently with its own snapshot rollback mechanism, enabling parallel task execution without cross-contamination. This isolation strategy is equally effective for shortcut actions, ensuring that they can be executed concurrently without interference.
Real-World Side Effect Mitigation.
All benchmark tasks execute within our controlled test environment using dedicated test accounts. Actions such as adding an item to cart, or sending emails affect only these isolated test accounts and never impact real users or production systems. This design ensures ethical evaluation practices while maintaining task authenticity.
A.3 Details of Action Space
@@
A.4 Details of Predefined Shortcuts Knowledge Base
@@
Our shortcut knowledge base was created to provide agents with an authentic yet challenging environment. We collect and design shortcuts for 11 selected applications, ensuring their relevance to real-world usage patterns. Importantly, we focus on collecting shortcuts that represent commonly used functionalities of these applications, which tend to be stable across app versions.
  • API Collection: The APIs in our knowledge base are intent-accessible endpoints rather than private or non-exported APIs. We identify documented endpoints from official Android documentation and use static analysis of application packages to locate additional endpoints exposed through standard Android IPC. We retain only APIs that can be invoked by standard ADB commands, such as adb shell am start, with Intent actions, data URIs, and extras. They do not require root access, debug signatures, hidden permissions, or emulator system modifications.
@@
  • RPA Script Design: RPA scripts are designed for common, highly repetitive sub-tasks (e.g., adding an item with specific options to a shopping cart). These scripts are implemented using UI tree-based element identification, which provides better robustness against UI layout changes compared to coordinate-based approaches. The scripts can be executed as a single action, serving as a high-level shortcut. This design enables us to test an agent’s ability to utilize complex, pre-built automation routines.
Stability of Shortcuts.
The stability of shortcuts across application versions is a critical consideration for real-world deployment. Our collection strategy deliberately focuses on core, commonly used functionalities that applications typically maintain across updates. API and Deep-Link shortcuts exhibit high stability, as they interface with core application functions that rarely change across versions. These shortcuts are based on documented or intent-accessible endpoints and declared URI schemes, which developers typically maintain for backward compatibility. RPA scripts in our knowledge base are implemented using UI tree-based element identification rather than fixed coordinates. This approach leverages the structured view hierarchy, which remains more stable across application updates compared to pixel-based coordinate systems, resulting in enhanced robustness against UI layout changes.
To validate the robustness of our shortcut knowledge base, we have tested all shortcuts across at least five different versions of each application, spanning updates released over six months. Our validation confirms that all shortcuts in our knowledge base remain functional across these versions without modification, demonstrating their practical viability for real-world agent deployment.
@@
Human - - 9.272 1.000 - - - -
MobileAgentV2 Gemini-2.0-Flash 0.250 12.881 1.280 472.426 40.406 0 0
MAS-MobileAgent Gemini-2.0-Flash 0.402 9.087↑29% 0.719↑44% 335.181 39.597 3.565 0.505
MobileAgentV2 Gemini-2.5-Pro 0.446 12.424 1.058 1013.386 120.212 0 0
MAS-MobileAgent Gemini-2.5-Pro 0.641 9.109↑27% 0.613↑42% 682.547 99.780 1.348 0.345
@@
Human - - 17.660 1.000 - - - -
MobileAgentV2 Gemini-2.0-Flash 0 31.553 - 952.660 106.501 0 0
MAS-MobileAgent Gemini-2.0-Flash 0.234 24.132↑24% 0.938 403.798 71.350 5.362 0.709
MobileAgentV2 Gemini-2.5-Pro 0.170 32.617 1.247 2053.133 227.128 0 0
MAS-MobileAgent Gemini-2.5-Pro 0.617 20.404↑37% 0.829↑34% 1441.586 189.836 3.128 0.320
Table 5: Evaluation results of different base models. MS: Mean Steps; MSRS: Mean Step Ratio on Successful tasks; MET: Mean Execution Time (seconds); MToC: Mean Token Cost (in thousands); MSC: Mean Shortcut Call count; S2GR: Shortcut-to-GUI action Ratio.
A.5 Details of Tasks in MAS-Bench
@@
This suite comprises 139 tasks strategically designed across 11 popular real-world applications. The tasks are categorized into 92 single-app tasks and 47 cross-app tasks, with their difficulty distribution illustrated in Table 7. We have also curated a corresponding knowledge base of predefined shortcuts to support these tasks. Table 17 presents a breakdown of the selected applications and the number of single-app tasks and shortcuts associated with each.
Table 6 summarizes the task coverage and representative examples. Single-app tasks cover shopping, local services, information browsing, media, productivity, health, and navigation workflows. Approximately 37% of single-app tasks are designed as incremental variations, where later tasks extend earlier ones with additional constraints or sub-goals. Cross-app tasks emphasize inter-app coordination and are mostly constructed by composing foundational single-app operations into information transfer, sharing, planning, and multi-app workflows.
Task Type Category # Tasks Representative Apps Example Task Pattern
Single-app Shopping and local services 28 Amazon, Booking.com, Yelp Search, filter, and verify products, hotels, attractions, or restaurants within one app.
Single-app Information and media browsing 26 BBC News, Chrome, YouTube Navigate news sections, search the web or videos, and save or play selected content.
Single-app Productivity, health, and navigation 38 Contacts, Fitbit, Gmail, Calendar, Maps Create contacts, draft emails, add calendar events, log health records, or obtain map directions.
Cross-app Information management 16 YouTube, Calendar, Booking.com, Contacts, BBC News Retrieve information in one app and record, schedule, or reuse it in another app.
Cross-app Web shopping and travel planning 11 Amazon, Booking.com, Gmail, Calendar, Chrome Search products, hotels, flights, or rentals and transfer the result to email, calendar, or another app.
Cross-app Multi-app coordination 11 Maps, Yelp, Gmail, Calendar, BBC News Coordinate several apps to complete chained goals such as finding places and sharing details.
Cross-app Social sharing 5 Maps, BBC News, Gmail Find content such as locations, photos, or articles and share the result through email.
Cross-app Media and entertainment 4 YouTube, Chrome, Fitbit, Calendar Search or consume media content and connect it with logging or scheduling actions.
Table 6: Task categories and representative examples in MAS-Bench. MAS-Bench contains 92 single-app tasks and 47 cross-app tasks across 11 real-world Android applications.
Human Baseline Collection Methodology.
To establish optimal step counts for efficiency metrics (MSR and MSRS), we conducted a systematic human expert annotation process. Three experienced mobile device users were recruited to independently complete each of the 139 tasks in MAS-Bench. Each expert was instructed to complete tasks as efficiently as possible using only pure GUI operations, specifically standard interactions such as taps, swipes, and text input, without access to any programmatic shortcuts. For each task, we recorded the number of steps taken by each expert and selected the minimum step count across the three demonstrations as the human baseline. This methodology ensures that our baseline represents the shortest achievable pure-GUI path under realistic conditions, providing a fair and reproducible reference for agent efficiency evaluation. The resulting human baselines average 9.27 steps for single-app tasks and 17.66 steps for cross-app tasks, reflecting the inherent complexity difference between these task categories.
@@
A.6 Evaluation Metrics
To comprehensively evaluate agent performance on MAS-Bench, we employ a multi-dimensional evaluation framework that assesses success, efficiency, and cost. Table 8 provides a complete summary of all metrics. Our evaluation framework is designed to encourage agents to efficiently complete tasks by intelligently utilizing existing and self-generated shortcuts, thereby achieving lower costs and greater efficiency.
Success Rate (SR).
Efficiency.
  • Mean Step Ratio (MSR): The ratio of agent steps to human steps (pure GUI operations). Lower values indicate better efficiency, with values below 1.0 representing fewer steps than the human.
  • Mean Step Ratio on Successful tasks (MSRS): The same ratio as MSR, computed only over successfully completed tasks.
Cost and Resource Utilization.
  • Mean Token Cost (MToC): The average number of tokens (in thousands) consumed by the LLM during task execution.
  • Shortcut to GUI Action Ratio (S2GR): This metric reflects an agent’s operational strategy preference by calculating the ratio of shortcut operations to GUI operations. A higher ratio indicates that the agent relies more heavily on efficient shortcuts rather than manual GUI actions.
Metric Symbol Definition Unit Dir.
Success
Success Rate SR Percentage of successfully completed tasks % $\uparrow$
Efficiency
Mean Steps MS Average number of steps (GUI actions + shortcuts) per task steps $\downarrow$
Mean Step Ratio MSR Ratio of agent steps to human baseline steps (pure GUI) ratio $\downarrow$
Mean Step Ratio on Successful MSRS MSR calculated only on successfully completed tasks ratio $\downarrow$
Mean Execution Time MET Average wall-clock time to complete a task seconds $\downarrow$
Cost and Resource Utilization
Mean Token Cost MToC Average LLM tokens consumed per task thousands $\downarrow$
Mean Shortcut Call Count MSC Average number of shortcut invocations per task count $\uparrow$
Shortcut Success Rate SSR Ratio of successful shortcut calls to total calls % $\uparrow$
Shortcut-to-GUI Ratio S2GR Ratio of shortcut actions to GUI actions ratio $\uparrow$
Table 8: Summary of evaluation metrics in MAS-Bench. Dir.: Direction of improvement ($\uparrow$ higher is better, $\downarrow$ lower is better). MSR and MSRS use human expert demonstrations (pure GUI operations) as the baseline. Values below 1.0 indicate agent performance exceeds human efficiency through shortcut utilization.
A.7 Task Difficulty Classification
MAS-Bench tasks are categorized into three difficulty levels to enable systematic evaluation across varying complexity. The classification criteria differ between single-app and cross-app tasks, reflecting their distinct operational characteristics.
Single-App Tasks.
For the 92 single-app tasks, difficulty is primarily determined by operational complexity measured through the number of required steps and the sophistication of sub-goals:
  • Level 1 (Easy): Tasks requiring $\leq 7$ steps. These involve basic, direct operations with a single primary goal, such as simple searches or navigation (e.g., searching for a product, opening a specific page).
  • Level 2 (Medium): Tasks requiring 8-15 steps. These involve multi-step workflows with intermediate complexity, including filtering operations, state verification, or coordination of multiple sub-goals (e.g., searching with filters and adding items to cart).
  • Level 3 (Hard): Tasks requiring $\geq 15$ steps. These involve complex workflows with multiple interdependent sub-goals, advanced filtering, cross-page navigation, and explicit state verification (e.g., multi-criteria product comparison with cart management and confirmation).
Cross-App Tasks.
For the 47 cross-app tasks, difficulty is determined by both the number of applications involved and the complexity of inter-app coordination:
  • Level 1 (Easy): Tasks involving 2 applications with straightforward information transfer, typically requiring $\leq 12$ steps. These tasks involve basic data lookup in one app and simple recording or sharing in another.
  • Level 2 (Medium): Tasks involving 2-3 applications with moderate coordination complexity, typically requiring 13-20 steps. These tasks involve filtering, processing, or transforming information across applications.
  • Level 3 (Hard): Tasks involving 3-4 applications with complex multi-app workflows, typically requiring $\geq 20$ steps. These tasks demand sophisticated inter-app dependencies, multi-stage data processing, and coordination across diverse application domains.
@@ Appendix B Details of Predefined Shortcut Evaluation
B.1 Detailed Results on Single-app and Cross-app Tasks
This section provides detailed performance for representative agents on single-app and cross-app tasks in MAS-Bench. Table 9 shows the results for 92 single-app tasks, while Table 10 presents the results for 47 cross-app tasks. These detailed results complement the overall weighted average results shown in Table 2 in the main text.
[…]
$S$
VH
[…]
Human - 9.272 1.000 - - - -
Agentic Workflow (Gemini-2.5-Pro)
M3A Rawles et al. (2024) 0.565 11.772 1.064 192.775 155.281 0 0
MobileAgent-E Wang et al. (2025d) 0.359 4.080 0.818 459.574 88.772 0.378 0.081
T3A Rawles et al. (2024) 0.511 11.750 1.056 137.641 346.382 0 0
   + MAS-T3A (Ours) 0.576↑13% 10.595↑10% 0.915↑13% 129.279↑6% 291.391↑16% 1.043 0.117
MobileAgentV2 Wang et al. (2024a) 0.446 12.424 1.058 1013.386 120.212 0 0
   + MAS-MobileAgent (Ours) 0.641↑44% 9.109↑27% 0.613↑42% 682.547↑33% 99.780↑17% 1.348 0.345
General-Purpose Models
Qwen2.5-VL-3B Bai et al. (2025b) 0.054 17.380 1.341 99.564 - 0 0
Qwen2.5-VL-7B Bai et al. (2025b) 0.022 17.315 1.625 167.892 - 0 0
Qwen3-VL-4B Bai et al. (2025a) 0.290 13.739 1.274 143.098 - 0 0
   + MAS-Qwen3-VL-4B (Ours) 0.315↑8.6% 15.011-9.3% 0.888↑30.3% 147.230-2.9% - 2.620 0.219
Qwen3-VL-8B Bai et al. (2025a) 0.326 12.761 1.179 148.021 - 0 0
   + MAS-Qwen3-VL-8B (Ours) 0.522↑60% 11.065↑13% 0.792↑33% 124.370↑16% - 1.239 0.129
Qwen3-VL-32B Bai et al. (2025a) 0.402 12.793 1.150 293.941 - 0 0
   + MAS-Qwen3-VL-32B (Ours) 0.543↑35.1% 11.859↑7.3% 0.810↑29.6% 269.291↑8.4% - 2.272 0.271
Qwen3-VL-235B Bai et al. (2025a) 0.478 11.989 1.113 166.083 - 0 0
   + MAS-Qwen3-VL-235B (Ours) 0.598↑25.1% 10.924↑8.9% 0.849↑23.7% 134.274↑19.2% - 1.989 0.219
GLM-4.5V Hong et al. (2025) 0.533 13.739 1.209 214.544 - 0 0
   + MAS-GLM-4.5V (Ours) 0.739↑38.6% 10.337↑24.8% 0.823↑31.9% 192.489↑10.3% - 0.957 0.110
Specialized GUI Models
UI-TARS-1.5-7B Seed (2025) 0.380 14.543 1.203 147.966 - 0 0
GUI-Owl-7B Ye et al. (2025) 0.402 11.554 1.279 132.698 - 0 0
ScaleCUA-7B Liu et al. (2025c) 0.163 13.554 1.202 98.037 - 0 0
   + MAS-ScaleCUA-7B (Ours) 0.174↑6.7% 14.685-8.3% 1.358-13.0% 111.684-13.9% - 0.011 0.001
ScaleCUA-32B Liu et al. (2025c) 0.283 14.065 1.273 113.134 - 0 0
   + MAS-ScaleCUA-32B (Ours) 0.272-3.9% 14.230-1.2% 1.245↑2.2% 115.300-1.9% - 0 0
MAI-UI-8B Zhou et al. (2025) 0.565 13.793 1.213 137.677 - 0 0
   + MAS-MAI-UI-8B (Ours) 0.652↑15.4% 13.065↑5.3% 1.130↑6.8% 131.573↑4.4% - 0.315 0.024
Table 9: Detailed performance on single-app tasks (92 tasks). Bold and underlined values denote the best and second-best results within each block, respectively. Methods marked with “+” are shortcut-augmented variants of their corresponding baselines. $S$: Screenshot; VH: View Hierarchy; MS: Mean Steps; MSRS: Mean Step Ratio on Successful tasks; MET: Mean Execution Time (seconds); MToC: Mean Token Cost (in thousands); MSC: Mean Shortcut Call count; S2GR: Shortcut-to-GUI action Ratio.
[…]
$S$
VH
[…]
Human - 17.660 1.000 - - - -
Agentic Workflow (Gemini-2.5-Pro)
M3A Rawles et al. (2024) 0.383 26.213 1.262 410.828 289.040 0 0
MobileAgent-E Wang et al. (2025d) 0.064 6.234 0.934 468.630 85.954 2.250 0.177
T3A Rawles et al. (2024) 0.340 23.596 1.087 257.015 625.970 0 0
   + MAS-T3A (Ours) 0.511↑50% 17.021↑28% 0.643↑41% 186.139↑28% 439.296↑30% 2.213 0.185
MobileAgentV2 Wang et al. (2024a) 0.170 32.617 1.247 2053.204 227.357 0 0
   + MAS-MobileAgent (Ours) 0.617↑263% 20.404↑37% 0.829↑34% 1479.197↑28% 192.052↑16% 3.122 0.333
General-Purpose Models
Qwen2.5-VL-3B Bai et al. (2025b) 0 29.936 - 162.886 - 0 0
Qwen2.5-VL-7B Bai et al. (2025b) 0.021 30.681 1.250 170.672 - 0 0
Qwen3-VL-4B Bai et al. (2025a) 0.106 25.000 1.272 249.891 - 0 0
   + MAS-Qwen3-VL-4B (Ours) 0.085-19.7% 30.957-23.8% 1.133↑10.9% 268.035-7.3% - 2.149 0.075
Qwen3-VL-8B Bai et al. (2025a) 0.128 24.043 1.192 240.755 - 0 0
   + MAS-Qwen3-VL-8B (Ours) 0.234↑83% 22.468↑7% 1.016↑15% 217.846↑9% - 1.447 0.073
Qwen3-VL-32B Bai et al. (2025a) 0.213 27.702 1.318 590.161 - 0 0
   + MAS-Qwen3-VL-32B (Ours) 0.255↑19.7% 23.849↑13.9% 0.923↑30.0% 532.476↑9.8% - 5.064 0.287
Qwen3-VL-235B Bai et al. (2025a) 0.298 25.638 1.106 222.624 - 0 0
   + MAS-Qwen3-VL-235B (Ours) 0.383↑28.5% 21.276↑17.0% 0.659↑40.4% 188.626↑15.3% - 4.340 0.261
GLM-4.5V Hong et al. (2025) 0.511 24.085 1.165 413.829 - 0 0
   + MAS-GLM-4.5V (Ours) 0.574↑12.3% 21.319↑11.5% 1.051↑9.8% 328.798↑20.5% - 1.234 0.071
Specialized GUI Models
UI-TARS-1.5-7B Seed (2025) 0.106 28.340 1.169 266.787 - 0 0
GUI-Owl-7B Ye et al. (2025) 0.085 23.426 1.161 237.539 - 0 0
ScaleCUA-7B Liu et al. (2025c) 0 27.277 - 171.917 - 0 0
   + MAS-ScaleCUA-7B (Ours) 0 28.766-5.5% - 199.968-16.3% - 0 0
ScaleCUA-32B Liu et al. (2025c) 0.128 29.277 1.310 197.481 - 0 0
   + MAS-ScaleCUA-32B (Ours) 0.106-17.2% 28.145↑3.9% 1.290↑1.5% 188.723↑4.4% - 0 0
MAI-UI-8B Zhou et al. (2025) 0.340 29.894 1.313 264.849 - 0 0
   + MAS-MAI-UI-8B (Ours) 0.447↑31.5% 27.851↑6.8% 1.274↑3.0% 243.632↑8.0% - 0.191 0.007
Table 10: Detailed performance on cross-app tasks (47 tasks). Bold and underlined values denote the best and second-best results within each block, respectively. Methods marked with “+” are shortcut-augmented variants of their corresponding baselines. $S$: Screenshot; VH: View Hierarchy; MS: Mean Steps; MSRS: Mean Step Ratio on Successful tasks; MET: Mean Execution Time (seconds); MToC: Mean Token Cost (in thousands); MSC: Mean Shortcut Call count; S2GR: Shortcut-to-GUI action Ratio.
B.1.1 Detailed Results of Single-app and Cross-app Tasks
The performance comparison between single-app and cross-app tasks reveals several critical insights into agent capabilities and the effectiveness of shortcut augmentation. Table 9 and Table 10 show the detailed results for representative agents.
Task Complexity and Performance Degradation.
Cross-app tasks demonstrate significantly higher complexity than single-app tasks, as evidenced by the substantial performance degradation across many agents. For instance, the baseline T3A achieves an SR of 51.1% on single-app tasks but drops to 34.0% on cross-app tasks, representing a 33% relative decrease. Similarly, MobileAgentV2 shows an even more pronounced decline from 44.6% to 17.0% (62% relative decrease). This trend is also observed in several Agent-as-a-Model baselines, with GUI-Owl-7B declining from 40.2% to 8.5%. The human baseline steps also increase from 9.272 to 17.660, confirming the inherent complexity difference between task types.
Amplified Benefits of Shortcuts on Cross-app Tasks.
Predefined shortcuts demonstrate disproportionately larger benefits on cross-app tasks compared to single-app tasks. MAS-MobileAgent improves MobileAgentV2’s SR by 44% on single-app tasks but achieves a remarkable 263% improvement on cross-app tasks (17.0% → 61.7%). Similarly, MAS-T3A shows a 13% improvement on single-app tasks versus 50% on cross-app tasks. This amplified effect stems from shortcuts’ ability to bridge cross-application boundaries. Complex inter-app transitions that are particularly challenging for GUI-only agents become single atomic actions with appropriate shortcuts.
Efficiency Gains Across Task Types.
The Mean Step Ratio on Successful tasks (MSRS) metric reveals consistent efficiency improvements from shortcut augmentation across both task types. For agentic workflows, MAS-MobileAgent achieves an MSRS of 0.613 on single-app tasks and 0.829 on cross-app tasks, substantially lower than the baseline MobileAgentV2’s 1.058 and 1.247, respectively. This indicates that shortcut-augmented agents consistently execute paths closer to optimal solutions. Notably, MobileAgent-E, despite its low MS due to frequent premature termination from errors, maintains a competitive MSRS of 0.818 on single-app tasks, demonstrating that when it does succeed, its self-generated shortcuts enable efficient execution.
Agent-as-a-Model Performance Characteristics.
The expanded Agent-as-a-Model results show a clear difference between general-purpose models and specialized GUI models. General-purpose models adapt more effectively to the hybrid action space: MAS-GLM-4.5V achieves the highest SR on both single-app (73.9%) and cross-app tasks (57.4%), while MAS-Qwen3-VL-32B and MAS-Qwen3-VL-235B obtain stable efficiency gains. In contrast, specialized GUI models do not automatically benefit from shortcuts. MAI-UI-8B improves with MAS, but ScaleCUA variants rarely invoke shortcuts and show marginal or negative changes. These results suggest that shortcut augmentation requires not only GUI perception but also reliable shortcut selection and tool-use reasoning.
Cost and Efficiency.
Token cost analysis reveals that shortcut augmentation provides substantial efficiency gains while marginally increasing cost. On single-app tasks, MAS-MobileAgent reduces token cost by 17% compared to MobileAgentV2, while on cross-app tasks, the reduction is 16%. Mean execution time (MET) improvements are particularly pronounced on cross-app tasks: MAS-T3A reduces MET by 28% compared to T3A (257s → 186s), demonstrating that shortcuts not only improve success rates but also significantly accelerate task completion through reduced step counts and more efficient execution paths.
B.2 Details of MAS-MobileAgent and MAS-T3A